OpenStudy (anonymous):

verify the identity cos(2A)=1-2sin^A

OpenStudy (gorv):

u missed something in Q

OpenStudy (camper4834):

verified

OpenStudy (gorv):

\[\cos2A=1-2\sin^2A\]

OpenStudy (gorv):

\[\sin^2A=1-\cos^2A\]

OpenStudy (anonymous):

yeah that's what i meant.

OpenStudy (gorv):

\[\cos2A=1-2*(1-\cos^2A)\]

OpenStudy (gorv):

what u got after solving this ??

OpenStudy (gorv):

=1-2+2cos^2A =-1+2cos^2A